OpenStudy (anonymous):

HELP WITH 1 MATH QUESTION PLZ!

OpenStudy (bloomlocke367):

I'll try. what's your question?

OpenStudy (anonymous):

its in the attachment

OpenStudy (bloomlocke367):

Oh, I've helped with a question like that before XD okay... do you know how to find the answer?

OpenStudy (bloomlocke367):

@Pamela16

OpenStudy (anonymous):

oh ok know i dont and i really need it done

OpenStudy (bloomlocke367):

okay.. well what you should know is that STUV and WXYZ are corresponding to each other, meaning ST corresponds WX and UV corresponds with YZ.. do you understand what I mean?

OpenStudy (bloomlocke367):

it gives you the length of ST and asks for WX with a ratio of 3:1... so, you could just write a proportion.

OpenStudy (bloomlocke367):

\[\frac{ ST }{ 3 }=\frac{ WX }{ 1 }\]\[\frac{ 39mm }{ 3 }=\frac{ ? }{ 1 }\] so now, just cross multiply and divide. what do you get?

OpenStudy (anonymous):

i dont get it :\

OpenStudy (bloomlocke367):

what don't you get?

OpenStudy (bloomlocke367):

I'm here to help.